how can you find a slope? using the coordinates (4,1) and (2,5)?
Answer by Alan3354(69443)

how can you find a slope? using the coordinates (4,1) and (2,5)?
---------
Slope is the change in y divided by the change in x.
y: 1-5 = -4
x: 4-2 = 2
slope, m = -4/2 = -2
--------
Sometimes it's called rise over run.